OpenStudy (anonymous):

3(2k-2)=2(4k-11) how would you show work and solve?

OpenStudy (diyadiya):

6k-6=8k-22 22-6=2k 16=2k k=8

OpenStudy (anonymous):

6k-6=8k-22 22-6=2k 16=2k k=8

OpenStudy (anonymous):

lol i did same thing

OpenStudy (diyadiya):

copyied xD

OpenStudy (anonymous):

guess i don't need to answer then LOL

OpenStudy (diyadiya):

as you wish !

OpenStudy (anonymous):

actually its -2 in front of the second set of parenthesses. sorry.

OpenStudy (anonymous):

In that case... k=2!!! Solve the same way, though. Distributive property, and all that blah, blah, blah.

OpenStudy (diyadiya):

ROFL !!@c108...

OpenStudy (anonymous):

i distributed but am not sure after that.

OpenStudy (anonymous):

3(2k-2)=-2(4k-11) 6k-6=-8k+22 14k=28 k=2 and we are done :)